OpenStudy (anonymous):

LaToya decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a beginner's fee of $100 and a monthly fee of $38. If x represents the number of months that LaToya is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time. 100 + 38x = C LaToya has allocated a maximum of $328 to spend on her gym membership. Which number line shows the possible number of months that LaToya can be a member of the gym?

OpenStudy (joehoffy):

this is an easy one. 100+38x=404 100-100+38x=404-100 38x=303 38/38x=304/38 x=8

OpenStudy (joehoffy):

can i get a medal? its the blue button that says best response. click it. thanks!